OpenStudy (anonymous):

What is the rate of change for f(x) = 4 sin x − 2 on the interval from x = 0 to x = pi over 2?

OpenStudy (calculusfunctions):

Rate of over an interval is the same as the average rate of change (slope of a secant). The average rate of change of a function f(x) over the interval\[x _{1} \le x \le x _{2}\]is\[\frac{ f(x _{2})-f(x _{1}) }{ x _{2}-x _{1} }\]